Spectrophotometric properties of a particular indi
Answer #1

A = e*C*L

4.97*10^4 *C*1 = 0.148

C = 2.977*10^-6

These concentration is the concentration after adding to the solution.

C1V1 = C2V2

2.977*10^-6*40 = x*20

x = 5.954*10^-6 M is the molarity of KOH solution